Solve the following systems algebraically.
{ y = -x + 5
{ y = x + 3

Found 2 solutions by richwmiller, stanbon:
Answer by richwmiller(17219)

add them
2y=8
y=4
4=x+3
x=1

Answer by stanbon(75887)

Solve the following systems algebraically.
{ y = -x + 5
{ y = x + 3
----
Add each column to get:
y+y = -x+x + 5+3
-----
2y = 0 + 8
2y = 8
Divide both sides by 4::
y = 4
----
Solve for "x":
y = x + 3
Substitute 4 for y:
4 = x + 3
Subtract 3 from each side to get::
x = 1
------------------
Cheers,
Stan H.
===================